# Fuss-Free Network: A Simplified and Efficient Neural Network for Crowd Counting

### v1

@@ Abstract
In the field of crowd-counting research, many recent deep learning based methods have demonstrated robust capabilities for accurately estimating crowd sizes. However, the enhancement in their performance often arises from an increase in the complexity of the model structure. This paper introduces the Fuss-Free Network (FFNet), a crowd counting deep learning model that is characterized by its simplicity and efficiency in terms of its structure. The model comprises only a backbone of a neural network and a multi-scale feature fusion structure. The multi-scale feature fusion structure is a simple architecture consisting of three branches, each only equipped with a focus transition module, and combines the features from these branches through the concatenation operation. Our proposed crowd counting model is trained and evaluated on four widely used public datasets, and it achieves accuracy that is comparable to that of existing complex models. The experimental results further indicate that excellent performance in crowd counting tasks can also be achieved by utilizing a simple, low-parameter, and computationally efficient neural network structure.
keywords:
crowd counting , Fuss-Free Network , multi-scale feature fusion , focus transition module , simplified structure
journal: Elsevier
[…]
[inst1]organization=School of Opto-Electronic and Communication Engineering, Xiamen University of Technology, city=Xiamen, country=China
1 Introduction
With the acceleration of urbanization and the increase in population mobility, effective monitoring and management of population density have become increasingly important. Rational regulation of population density not only concerns public safety but also affects social stability and the optimization of resource allocation [1]. Crowd counting technology, which has been receiving growing attention in the field of computer vision [2], has the potential to monitor and analyze crowd density in real-time, thereby helping to prevent safety incidents resulting from overcrowding. Moreover, it can provide essential data support for urban management, facilitating the optimization of resource allocation in crowded areas. Additionally, during the pandemic, this technology has the ability to monitor and regulate crowd movements, contribute to the development of public health strategies, and mitigate the transmission of the virus.
@@ 1 Introduction
Crowd counting is a highly challenging computer vision task. Figure 1 shows three representative images for crowd counting, illustrating the following challenges in crowd counting tasks: (1) Variations in scale and density within an image: In general, crowds that are closer to the camera tend to have larger scales and lower densities, while those further away exhibit the opposite characteristics. (2) Object occlusion and overlap in high density crowds: High density crowd counting tasks frequently involve analyzing crowd images where a significant number of individuals are present, and situations such as object occlusion and overlap between individuals are commonly encountered. (3) The effects of lighting and background variations: The characteristics of individual features can vary due to differences in lighting conditions, which can include the creation of shadows. Moreover, the complexity and variability of the background environment can introduce interference factors, such as object occlusion. Moverover, crowd counting often involves the real-time processing of significant amounts of image or video data to fulfill the requirements of real-time crowd flow monitoring.
@@
To overcome these challenges, mainstream methods of crowd counting rely on the powerful performance of deep learning models. These models are designed and trained to generate crowd density images at the pixel level. Subsequently, an accurate estimation of the total number of individuals is achieved by using a pixel-level crowd density image. To address the challenge of scale and density variations present in images, numerous methods have employed a multi-branch structure. This approach allows for the utilization of multi-scale features, effectively tackling the issue. Among them, some methods employ dilated convolutions with different dilation rates in a multi-branch structure, enabling the model to have receptive fields of different scales. Furthermore, to tackle challenges such as object occlusion, overlap in high density crowds, and the effects of lighting and background variations, almost all methods focus on enhancing the feature extraction capability of the model, primarily by selecting or designing various network architectures and employing different attention mechanisms. Optimizing the loss function yields significant effectiveness as well. For instance, in order to enhance model supervision, the loss function can be adjusted to ensure that the model establishes a more precise correspondence between the predicted density and the labeled location.
Different research works adopt different perspectives and consider various rules. Through intricate and sophisticated network structural designs and algorithm optimizations, existing deep learning models have made remarkable progress in improving the accuracy of crowd counting. However, the complexity of the model to a large extent hinders subsequent researchers from modifying and further improving the model. For example, efforts to simplify existing models with complex structure to reduce the consumption of computing resources may significantly impact the accuracy of crowd counting and the models’ adaptability in complex scenarios. However, for structurally simple models, it is relatively easier to modify the models while minimizing the performance decrease. Hence, the reduction in complexity of the model structure is beneficial for the subsequent advancement of crowd counting. In the pursuit of simple structure but also need to consider the counting performance, achieving a balance between simplifying structure and enhancing performance is a challenge that we need to pay attention to.
For the aforementioned reasons, our research objective is to design a crowd-counting model that strikes a balance between simplifying the model structure and improving counting performance for crowd counting. In terms of specific implementation, we first constructed a powerful backbone network, which is used to accurately capture the initial features of crowds from raw images. Next, we intricately designed a focus transition module, which efficiently integrates and refines key feature information in dynamic crowd scenes, ensuring the model’s effective adaptation to complex environmental changes. Finally, by implementing a practical and feasible multi-scale feature fusion strategy, we achieved deep integration and optimization processing of the extracted multi-level features. This ensures accurate identification and counting of crowds at different scales. Research shows that excellent crowd counting models can be designed with just a simple structure.
In summary, this paper’s contributions can be outlined in three aspects:
  1. 1.
@@
    We propose a crowd-counting neural network with a simple structure that consists of only a backbone and a multi-scale feature fusion structure, in contrast to other complex models. However, it achieves results comparable to the state-of-the-art, even with low parameters and computational requirements.
  2. 2.
@@ 2 Related Work
Crowd counting is a task of estimating the number of people in a specific location. Early research employed human detection methods [3, 4], where researchers used detectors to detect human in images and accumulated the count of each detected person to obtain the crowd count result in the image. The limitation of this method lies in the significant impact of occlusions between persons in crowded environments or densely packed crowds on the performance of detectors, leading to a decrease in counting accuracy. Subsequently, direct regression methods [5] were proposed, which count the crowd directly from images. However, this method often cannot provide detailed information about individual distribution and positioning, and it is sensitive to model assumptions and parameter selection, which can affect accuracy in complex or highly variable density scenes. Therefore, methods based on density map regression [6, 7] have become mainstream, where the final crowd count is obtained by integrating density map estimates from images. Relying on the powerful representation capability of convolutional neural networks (CNNs), the majority of methods in this field are built upon the framework of CNNs. There are also many methods leveraging the strong feature extraction ability of transformers. To further improve performance, researchers have developed and integrated various technological and architectural innovations. In this section, we will review the relevant work on crowd counting from three aspects: model design, loss function design, and algorithm strategy design.
2.1 Model Design for Crowd Counting
With the development of deep learning, various methods for model design in crowd counting have emerged. These methods include introducing multi-scale fusion strategies, integrating attention mechanisms, and employing transformers. Multi-scale feature fusion methods attempt to address the challenge of large variations in crowd scale or size diversity by integrating features from multiple different levels. Zhang et al. [6] adopt a multi-column CNN architecture, where each column CNN utilizes filters with different receptive fields to adapt to scale variations. Jiang et al. [9] adopt an encoder-decoder structure that integrates multi-level features through skip connections across layers, aggregating low-level spatial details and high-level semantic information in a bottom-up manner. This enables accurate prediction from the entire image input to high-resolution density maps. Liu et al. [10] propose a novel deep learning architecture that utilizes spatial pyramid pooling to extract context information at different scales. It then uses these scale-aware features to regress the final density map. Ma et al. [11] introduce a multi-scale mechanism that utilizes information provided by different layers in the encoder to generate feature maps with varying receptive field sizes. Zhang et al. [12] designed a scaled adaptive convolutional neural network called SaCNN. This network combines features from multiple levels to enable the model to flexibly correspond to scale and viewpoint differences in the input image. Han et al. [13] employ a feature selection and inheritance learning approach, along with mask selection and inheritance loss, to perform multi-scale fusion. These methods achieve good results by effectively integrating multi-scale features.
Attention mechanisms are also an efficient method. Cao et al. [14] proposed SANet, which utilizes multi-scale convolutions and an adaptive global attention mechanism to handle crowd images of different sizes and densities. Jiang et al. [15] adopt a dual-branch structure, where one branch generates adjusted intermediate density maps and scale factors. It then corrects the unevenness of crowd density in the image by combining it with attention masks generated by the other branch, ultimately producing refined attention-based density maps. Lin et al. [16] employ a learnable global attention module and a local partition attention module, achieving decent performance in counting accuracy.
@@ 2.1 Model Design for Crowd Counting
Given the remarkable success of transformer networks in the field of natural language processing (NLP), researchers have also begun to explore their application in crowd counting as a core feature extraction module. Sun et al. [17] utilize transformer models for feature extraction and information exchange by overlapping and concatenating image blocks and adding contextual tokens. Qian et al. [18] introduce a “U-shaped” multi-scale transformer network for crowd counting and introduce a new loss function to supervise regression, significantly improving counting accuracy.
In addition to the aforementioned methods, there are some other designs that contribute to crowd counting. The CSRNet proposed by Yan et al. [2] consists of a convolutional neural network serving as the frontend for 2D feature extraction, and it utilizes dilated convolutional neural networks as the backend to provide a larger receptive field and replace pooling operations. Similarly, Tian et al. [8], besides employing a transformer network that focuses more on spatial aspects, also utilize a regression head composed of dilated convolutions to expand the receptive field, achieving good counting results as well.
2.2 Loss function design for crowd counting
@@
Many studies have designed loss functions tailored to the characteristics of crowd counting. Ma et al. [20] no longer rely on low-quality density maps generated from Gaussian kernels but construct a density contribution probability model from point annotations, using Bayesian inference for loss estimation. Wang et al. [21] treat crowd counting as a distribution matching problem, employing Optimal Transport (OT) to find the best match between the generated density map and the point annotations. Yan et al. [22] proposed a multi-resolution loss function, PML, which can learn to better fit the posterior distribution using multi-scale information, thereby improving counting performance.
2.3 Algorithmic Strategies design for crowd counting
@@
In addition to model design and loss function design, there are also studies that approach crowd counting from algorithmic perspectives. Song et al. [23] developed a novel approach named P2PNet, which directly predicts head coordinates and their confidences in images and matches points to their ground-truth targets using the Hungarian algorithm. Lin et al. [19] decompose the pedestrian counting process into point-querying processes and use a custom quad-tree structure to adapt to crowds of different densities. Additionally, it employs a recursive rectangular window attention mechanism to improve computational efficiency.
3 Method
@@
Figure 2: This figure presents the architectural designs of four different crowd counting models, including SASNet, STEERER, CCTrans, and our simplified architecture.
@@
With the continuous evolution and development of deep learning technology, high-performance crowd counting models have achieved significant improvements in accuracy, but they also face the problem of increasing model complexity. As depicted in Figure 2, the three exemplary high-performance models (CCTrans, STEERER, and SASNet) demonstrate exceptional accuracy, but their underlying structures are highly complex. Specifically, SASNet utilizes multiple decoder blocks along with a complex fusion header design and interleaved fusion mechanisms, resulting in a complex architecture. STEERER relies on the intricate and computationally demanding HRNet48 as its backbone network and designs a complex fusion module. Although the CCTrans appears to be lightweight, it actually adds more complexity to the model overall because it is based on the massive parameter Twins-SVT pre-training network, even though it only attaches a regression header after multi-branch fusion.
Based on this, the core of our research is to conceive and build a more streamlined and efficient framework for crowd counting models. The fourth framework in Figure 2 is precisely the simplified solution we propose. This model adopts the compact yet powerful ConvNeXT-Tiny as the base backbone network, and it comprises a multi-scale feature fusion structure that includes a focus transition module. We refer to this model as Fuss-Free Network(FFNet).
FFNet is specifically illustrated in Figure 3. Firstly, the input image goes through the feature extraction backbone in the framework to extract features at three different scales. Then, these three features pass through the focus transition module to generate more semantically rich features. Next, each stage’s features go through channel concatenation and spatially-aware fusion via the framework’s multi-scale fusion. Finally, regression heads are used to regress the density maps. The final density maps and the integrated counts will be used to construct the loss function.
@@
Figure 3: The network architecture of FFNet. H, W, C denote the height, width, and number of channels of the features, respectively. Gap represents global average pooling, Mean denotes mean operation, Conv stands for convolution, $\times1$Conv denotes convolution with a $\times1$ kernel, Dyconv represents dynamic convolution, MLP refers to multilayer perceptron, Sigmoid is the activation function, and the circles with “×”, “+”, and “C” represent element-wise multiplication, element-wise addition, and channel concatenation operations, respectively.
3.1 Feature Extraction Backbone
Different from most crowd counting studies, we employ ConvNeXt-Tiny [24] as the feature extraction backbone for the multi-scale fusion framework. This network integrates key design principles of transformers, such as layer normalization and multi-head attention mechanisms, and implements them in the ConvNeXt block structure. This combination allows ConvNeXt-Tiny to not only maintain the powerful processing capability of CNNs for spatial local features but also significantly enhance the understanding and utilization of global information.
Compared to the deep structures of VGG [25] and the residual learning of ResNet [26], ConvNeXt-Tiny provides a more efficient computational path and a simplified network architecture. Its streamlined and efficient characteristics ensure that the model can achieve superior performance in visual tasks such as image recognition compared to traditional models. Therefore, it also aligns better with our desire for a concise and efficient network design. For crowd counting tasks, dealing with objects of various scales has always been a significant challenge. A widely used solution is to generate features at different resolutions. Thus, we will utilize the ConvNeXt-Tiny pretrained model to build a multi-stage hierarchical feature extraction structure, where each stage can capture feature representations at different levels and resolutions. This approach enables us to better address the multi-scale challenges of crowd counting while ensuring the simplicity of the framework.
3.2 Multi-scale Fusion Strategy
@@
For features at different levels of resolution, we need to align and fuse them. To avoid introducing too much additional computational cost and adhere to the principle of simplicity and efficiency, we propose three simple designs for multi-scale fusion. As shown in the figure 3, the “Concatenate” fusion directly applies transposed convolution to features $S_{3}$ and $S_{2}$ for upsampling to match the spatial size of $S_{1}$, and then fuses the features through channel concatenation. The “Addition” fusion first reduces the number of channels in features $S_{3}$ and $S_{2}$ to match the channel count of $S_{1}$. Before performing element-wise addition, transposed convolution operations are used to align the spatial size of the features. The “Stepwise Addition” fusion involves progressively upsampling and combining features of different scales, similar to the design of FPN [28]. In detail, these designs employ transpose convolution instead of interpolation operations because transpose convolution can better preserve and reconstruct the spatial structure information and feature hierarchy of the image. Compared to simple methods like bilinear interpolation for upsampling, transpose convolution enhances the output resolution by learning weight parameters, allowing for a more flexible and adaptive generation of high-quality feature maps. This is particularly significant for reconstructing complex crowd distribution scenes. The above three designs performed well in experiments, with the “Concatenate” fusion achieving the best results. We default to this design in the FFNet framework.
3.3 Focus Transition Module
Directly fusing features from different levels and resolutions of the backbone may lead to information redundancy due to misalignment in spatial dimensions and an excessive number of channels. This could potentially result in adverse effects such as decreased learning efficiency and inaccurate information propagation. Especially in crowd counting tasks, directly fusing features may fail to accurately retain spatial details from lower layers and lack an understanding of the global scene layout. This can lead to biases in estimating dense or sparse crowd distributions. Therefore, to address this issue and fully exploit the potential of the framework, we designed a focus transition module. It effectively focuses on dynamic and static features, achieving efficient dimensionality reduction and feature extraction to facilitate the transition.
Inspired by [27], we employed a novel dynamic convolution in the design stage, as illustrated in Figure 5.
Figure 5: Dyconv. FC stands for fully connected layer, Gap represents global average pooling, Conv denotes convolution, Sigmoid and Relu are activation functions, multiply refers to element-wise multiplication, $A_{nk}$, $A_{sk}$, $A_{ik}$, $A_{ok}$ represent attention along the dimensions of the number of convolution kernels, convolution kernel size, input channels, and output channels, respectively, and $W_{k}$ represents the weights of different convolution kernels.
Its core idea is to dynamically generate convolution kernel weights during the model runtime based on the input data and the network’s learned state, rather than using predefined and fixed weight matrices. Dynamic convolution will flexibly capture features across the four key dimensions: input channels, output channels, kernel size, and the number of kernels within the convolution. It utilizes average pooling and convolution for attention and then dynamically weights them. It can adaptively adjust the parameters of the convolutional kernel to better match the features of the input data. This adaptive capability enables dynamic convolution to effectively extract features from data of different scales and complexities, thereby enhancing the model’s ability to handle diverse inputs. In terms of algorithms, we assume X to be the input of dynamic convolution. The process of dynamic convolution attention extraction and dynamic weighting can be defined as:
where A(X) represents the attention calculation, AvgPool denotes average pooling, FC stands for fully connected layer, Conv indicates convolution operation, n represents the number of convolutional kernels, and $\sigma$ is the sigmoid activation function; Y(X) represents the dynamic convolution weighting process, $A_{nk}$, $A_{sk}$, $A_{ik}$, $A_{ok}$ are the attention weights on four dimensions: the number of convolution kernels, kernel size, input channels, and output channels, respectively, $W_{k}$ represents the weights of different convolution kernels, and C denotes the input optimized by weights.
In terms of specific design, as illustrated in Figure 3, the entire structure combines feature weight optimization, dynamic perception, and residual components. The feature output from the backbone is first subjected to adaptive feature weight optimization. This mechanism strengthens and filters features along the channel dimension, ensuring that the model focuses on the features crucial for crowd statistics. Then, the stacked dynamic convolutions automatically adjust filter parameters for different inputs to capture dynamic and static feature changes in the scene more effectively, achieving feature extraction in the process. Finally, the module performs spatial feature weight optimization, allocating different attention to each pixel in the image, enabling the model to automatically focus on key areas of crowd distribution while ignoring irrelevant background noise. In addition, we embed residual structures throughout the entire process to ensure that the original low-level feature details are preserved and integrated with the higher-level abstract features refined through multiple convolutional layers. It is worth noting that while the focus transition module enhances key features, it also implicitly achieves feature dimensionality reduction. This dimensionality compression and optimization, while preserving the integrity of key information, help reduce the computational complexity of the model, thereby improving operational efficiency. In detail, after each dynamic convolution, there is a ReLU activation layer and a BatchNorm layer to ensure the non-linear expression capability of features and the stability of normalization. The entire process of the focus transition can be represented by the following formula:

### v3

# The Effectiveness of a Simplified Model Structure for Crowd Counting

@@ Abstract
In the field of crowd counting research, many recent deep learning based methods have demonstrated robust capabilities for accurately estimating crowd sizes. However, the enhancement in their performance often arises from an increase in the complexity of the model structure. This paper discusses how to construct high-performance crowd counting models using only simple structures. We proposes the Fuss-Free Network (FFNet) that is characterized by its simple and efficieny structure, consisting of only a backbone network and a multi-scale feature fusion structure. The multi-scale feature fusion structure is a simple structure consisting of three branches, each only equipped with a focus transition module, and combines the features from these branches through the concatenation operation. Our proposed crowd counting model is trained and evaluated on four widely used public datasets, and it achieves accuracy that is comparable to that of existing complex models. Furthermore, we conduct a comprehensive evaluation by replacing the existing backbones of various models such as FFNet and CCTrans with different networks, including MobileNet-v3, ConvNeXt-Tiny, and Swin-Transformer-Small. The experimental results further indicate that excellent crowd counting performance can be achieved with the simplied structure proposed by us.
Index Terms:
crowd counting, simplified structure, Fuss-Free Network, multi-scale feature fusion, focus transition module.
I Introduction
Crowd counting, which has been receiving growing attention in the field of computer vision [1], has the potential to monitor and analyze crowd density in real-time, thereby helping to prevent safety incidents resulting from overcrowding. Moreover, it can provide essential data support for urban management, facilitating the optimization of resource allocation in crowded areas. Additionally, during the pandemic, it has the ability to monitor and regulate crowd movements, contribute to the development of public health strategies, and mitigate the transmission of the virus.
@@ I Introduction
Crowd counting is a highly challenging computer vision task. Figure 1 shows three representative images for crowd counting, illustrating the following challenges in crowd counting tasks: (1) Variations in scale and density within an image: In general, crowds that are closer to the camera tend to have larger scales and lower densities, while those further away exhibit the opposite characteristics. (2) Overlap occlusion and object in high density crowds: High density crowd counting tasks frequently involve analyzing crowd images where a significant number of individuals are present, and situations such as object occlusion and overlap between individuals are commonly encountered. (3) Lighting and background variations: The characteristics of individual features can vary due to differences in lighting conditions, which can include the creation of shadows. Moreover, the complexity and variability of the background environment can introduce interference factors, such as object occlusion.
@@
To overcome these challenges, mainstream methods of crowd counting rely on the powerful performance of deep learning models. These models are designed and trained to generate crowd density images at the pixel level. Subsequently, an accurate estimation of the total number of individuals is achieved by using a pixel-level crowd density image. In order to tackle the issue of scale and density variations in images, many approaches have adopted a multi-branch structure. Among them, some methods employ dilated convolutions with different dilation rates in different branchs, enabling the model to have receptive fields of different scales. Furthermore, to tackle challenges such as object occlusion, overlap in high density crowds, and lighting and background variations, most methods focus on enhancing the feature extraction capability of the model, primarily by selecting or designing powerful network architectures. Optimizing the loss function yields significant effectiveness as well. For instance, to enhance model supervision, the loss function can be adjusted to ensure that the model establishes a more precise correspondence between the predicted density and the labeled location.
In this paper, we focus on the impact of structural complexity on modifying models, discussing whether it is possible to find structurally simple models that achieve less performance decrease compared to complex models after modification. We strive to comprehensively consider the simplicity of the structure and the effectiveness of the model. Therefore, we opt for the ConvNeXT-Tiny architecture as our backbone, renowned for its simplicity and efficient design in convolutional neural networks. The multi-scale feature fusion structure we designed is very simple and allows us to effectively utilize information from different scales by fusing features from three different stages within the backbone. To simplify the structure, we only use concatenate operation on the features from three different branches. Moreover, we enhance the model’s effectiveness by including a carefully designed focus transition module before the fusion operation. In summary, this paper’s contributions can be outlined in three aspects:
  1. (1)
@@
    We present a crowd counting model with a simple structure that consists of only a backbone and a multi-scale feature fusion structure. Unlike other complex structure models, the fact that modifying our model, such as replacing different backbones, can achieve minimal performance loss demonstrates the effectiveness of this simple structure. Additionally, it achieves state-of-the-art (SOTA) results even with low parameters and computational requirements.
  2. (2)
@@
Early research employed human detection methods [2], where researchers used detectors to identify humans in an image and then counted each detected person. The drawback of this method is that occlusions between individuals in crowded or crowded environments greatly affect the performance of detectors, resulting in a decrease in counting accuracy. Direct regression methods [3] directly estimate the number of people. However, this approach often lacks detailed information about individual distribution and positioning, and it is sensitive to model assumptions, which may impact accuracy in complex or highly variable density scenarios. Nowadays, density map regression methods [4] have become popular, as they integrate density map estimates from images to determine the crowd count. These methods rely heavily on strong representation capabilities of networks, with most of them utilizing convolutional neural networks (CNNs). Additionally, some methods also leverage transformers. To enhance performance, researchers have made advancements in technology and architecture. In this section, we review the relevant research on crowd counting, focusing on three aspects: model design, loss function design, and other aspects of the design.
I-A Model Design for Crowd Counting
Most of the research focuses on the following three aspects: utilizing powerful backbones, designing multi-scale feature fusion strategies, and integrating attention mechanisms.
@@ I-A Model Design for Crowd Counting
Choosing the right backbone is the primary task for building a crowd counting model. Given the remarkable success of transformer networks in the field of natural language processing (NLP), researchers have also begun to explore their application in crowd counting as a core feature extraction module. Sun et al. [5] utilize transformer models for feature extraction and information exchange by overlapping and concatenating image blocks and adding contextual tokens. Qian et al. [6] introduce a “U-shaped” multi-scale transformer network for crowd counting and introduce a new loss function to supervise regression, significantly improving counting accuracy.
Multi-scale feature fusion methods attempt to address the challenge of large variations in crowd scale or size diversity by integrating features from multiple different levels. Zhang et al. [4] adopt a multi-column CNN architecture, where each column CNN utilizes filters with different receptive fields to adapt to scale variations. Liu et al. [7] propose a novel deep learning architecture that utilizes spatial pyramid pooling to extract context information at different scales. It then uses these scale-aware features to regress the final density map. Ma et al. [8] introduce a multi-scale mechanism that utilizes information provided by different layers in the encoder to generate feature maps with varying receptive field sizes. Han et al. [9] employ a feature selection and inheritance learning approach, along with mask selection and inheritance loss, to perform multi-scale fusion. Wang et al. [10] use tree-based enhancers to address variations in crowd scale and leverages multi-level auxiliaries to filter out background pixels while adapting to changes in crowd density. These methods achieve good results by effectively integrating multi-scale features.
Attention mechanisms are also an efficient method. Jiang et al. [11] adopt a dual-branch structure, where one branch generates adjusted intermediate density maps and scale factors. It then corrects the unevenness of crowd density in the image by combining it with attention masks generated by the other branch, ultimately producing refined attention-based density maps. Lin et al. [12] employ a learnable global attention module and a local partition attention module, achieving decent performance in counting accuracy. Wang et al. [13] adopted a multi-layer context fusion approach and introduced a directive attention mechanism to improve accuracy.
I-B Loss Function for Design Crowd Counting
@@
Some works have designed loss functions tailored to the characteristics of crowd counting tasks. Wang et al. [14] treated crowd counting as a distribution matching problem, employing Optimal Transport (OT) to find the best match between the generated density map and the point annotations. Yan et al. [15] proposed a multi-resolution loss function, progressive multi-resolution loss (PML), which can learn to better fit the posterior distribution using multi-scale information, thereby improving counting performance.
I-C Other Aspects
@@
In addition to model and loss function design, there are also some studies improving crowd counting methods from other different aspects. Song et al. [16] developed a novel approach named P2PNet, which directly predicts head coordinates and their confidences in images and matches points to their ground-truth targets using the Hungarian algorithm. Lin et al. [17] decompose the pedestrian counting process into point-querying processes and use a custom quad-tree structure to adapt to crowds of different densities. Additionally, it employs a recursive rectangular window attention mechanism to improve computational efficiency.
II Method
@@
Figure 2: This figure presents the architectural designs of four different crowd counting models, including SASNet, PET, CCTrans, and our simplified structure.
@@
In recent years, a series of high-performance crowd counting models based on deep learning have been proposed. The trade-off for achieving high performance in these models is an increase in the complexity of the model structure. Figure 2 illustrates a comparison between the network structures of three mainstream high-performance crowd counting models, namely CCTrans [18], PET [17], and SASNet [19], and our proposed model structure. More specifically, SASNet incorporates multiple decoder blocks, along with a sophisticated fusion header design and interleaved fusion mechanisms, resulting in a highly intricate architecture. PET employs the Quadtree structure to address the challenge of crowd counting and incorporates a complex Transformer-based encoder-decoder, substantially amplifying the model’s intricacy. Although the structure of CCTrans may seem simple, it actually introduces additional complexity to the overall model. This is because it is built upon the massive parameter Twins-SVT network, even though it only adds a regression header after multi-branch fusion.
The fourth structure in Figure 2 illustrates the simplified solution proposed by us. We utilize the compact yet powerful ConvNeXT-Tiny as the base backbone and integrate only a multi-scale feature fusion structure with a focus transition module. This model is referred to as Fuss-Free Network (FFNet). As depicted in Figure 3, FFNet comprises solely of a backbone and a simplified multi-scale feature fusion module. The backbone is responsible for extracting features at three different scales. The simplified multi-scale feature fusion module includes three branches, each equipped with a focus transition module. The features at three different scales are individually processed through their corresponding focus transition modules, resulting in the generation of more semantically rich features specifically related to crowded scenes. Finally, the features from the three branches are concatenated together, and then processed through a 1×1 convolution layer to merge the features into a single channel, resulting in the final density map. The source code for this project is available on GitHub at https://github.com/erdongsanshi/FFNet.
@@
Figure 3: The network architecture of FFNet. $H$, $W$, $C$ denote the height, width, and number of channels of the features, respectively. And the circles with “×”, “+”, and “C” represent element-wise multiplication, element-wise addition, and channel concatenation operations, respectively.
II-A Feature Extraction
Given our research objective of developing a crowd counting model with a balanced approach to simplifying the model structure, we opt for ConvNeXT-Tiny [20] as the backbone due to its notable attributes of simplicity and efficiency. ConvNeXts series build upon the foundation of a standard ResNet [21] while drawing inspiration from newer models proposed after 2020, such as vision Transformer and ResNeXt [22], with their innovative design and training techniques, for instance, one can apply techniques such as adjusting the stage compute ratio, utilizing Inverted Bottleneck and Large Kernel Sizes, and substituting ReLU with GELU, among others. By incorporating these operations, ConvNeXt-Tiny is able to preserve the strong spatial processing capabilities of CNNs for local features while also greatly improving the comprehension and utilization of global information.
ConvNeXt-Tiny is a smaller version of the original ConvNeXt model which provides a more efficient computational path and a simplified structure. Its streamlined and efficient design allows it to outperform traditional models in visual tasks like image recognition. This aligns with our goal of creating a concise and efficient network. In crowd counting tasks, dealing with objects of various sizes is a major challenge. One common approach is to generate features at different resolutions. To address this, To address this issue, we extract features from diverse stages of ConvNeXt-Tiny, thereby better adapting to the crowd counting task by capturing features at different hierarchical levels and resolutions from each stage.
II-B Multi-scale Feature Fusion
@@
The main objective of multi-scale feature fusion is to enhance model performance by effectively integrating features at various scales, which enables the model to capture both fine-grained details and overall contextual information simultaneously. As a result, the model’s accuracy, robustness, and generalization capabilities are improved, allowing it to better handle variations in object size, pose, appearance, etc.
@@
In order to adhere to the principle of simplicity and efficiency, we evaluate three commonly used and simple multi-scale feature fusion strategies and select the most efficient one. As shown in the Figure 3, the “Concatenate” fusion directly applies transposed convolution to features $S_{2}$ and $S_{3}$ for upsampling to match the spatial size of $S_{1}$, and then fuses the features through channel concatenation. The “Addition” fusion first reduces the number of channels in features $S_{3}$ and $S_{2}$ to match the channel count of $S_{1}$. Before performing element-wise addition, transposed convolution operations are used to align the spatial size of the features. The “Stepwise Addition” fusion involves progressively upsampling and combining features of different scales, similar to the design of FPN [23]. For the purpose of reconstructing complex crowd distribution features, we specifically select transpose convolution instead of interpolation operations, as it offers notable advantages such as learning capability, flexibility, spatial accuracy, and the ability to handle irregular patterns. This choice allows for more precise and adaptive upsampling, making it particularly impactful in the context of complex crowd distribution scenes.
After conducting experiments, we found that the above three fusion strategies showed satisfactory performance. Among them, the “Concatenate” fusion strategy yielded the best results. Therefore, we have chosen to adopt this design as the default fusion strategy in the FFNet framework.
II-C Focus Transition Module
The backbone was originally designed for classification tasks, so the features it outputs contain more general semantic information, which differs from the features required for crowd counting tasks. For example, the most crucial task in crowd counting is to extract features of the crowd. Although the backbone will adapt to crowd counting tasks during the training process, further processing after the three features at different scales will enhance the flexibility of the model, thus better adapting to crowd counting tasks. Furthermore, the direct fusion of the different scales features from backbone may result in information redundancy, potentially reducing learning efficiency. For the aforementioned reasons, we have developed a focus transition module that effectively focuses on dynamic and static features, enabling efficient dimensionality reduction and feature extraction to facilitate the transition.
According to the aforementioned requirements, we have devised the focus transition module (FTM) to transform the features of the three branches prior to fusion, while also adhering to principles of simplicity and efficiency. Figure 3 illustrates the main components of FTM, comprising dynamic convolution based [24] feature transformation, a channel attention mechanism, and a spatial attention mechanism. To begin with, we utilize a channel attention mechanism to capture the interdependencies among different channels of the feature map extracted from the backbone network. This mechanism assigns weights to each channel according to their significance, enabling the branch to concentrate on pertinent features while suppressing less informative ones. Subsequently, the dynamic convolution based feature transformation component is employed to enhance the feature map by effectively capturing both dynamic and static feature variations in the scene. Lastly, we implement a spatial attention mechanism to capture the interdependencies within the spatial of the feature map. This mechanism assigns weights to various spatial locations, allowing the branch to focus on spatial regions while suppressing irrelevant or noisy ones.
The most crucial component of FTM is the dynamic convolution based feature transformation, which includes dynamic convolution and residual connection. Dynamic convolution is employed in our model as an alternative to conventional convolution, effectively enhancing the model’s representational power without the need for increasing the depth or width of the network. Furthermore, we incorporate residual connections throughout the entire process to enhance gradient flow and facilitate the learning of intricate relationships. It is worth noting that while the FTM enhances key features, it also implicitly achieves feature dimensionality reduction. This dimensionality reduction, while preserving the integrity of key information, helpS reduce the computational complexity of the model. In detail, after each dynamic convolution, there is a ReLU activation layer and a BatchNorm layer to ensure the non-linear expression capability of features and the stability of normalization. The entire process of the FTM can be represented by the following formula: